Q:- A point moves so that the sum of the squares of its distances from the six faces of a cube given by x=±1, y=±1, z=±1 is 10 units . The locus of the point is


A. x^2+y^2+z^2=1


B. x^2+y^2+z^2=2


C. x+y+z =1


C. x+y+z=2

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Tapasranjan

planes are

x=±1, y=±1, z=±1

let the point is (x1,y1,z1)

 use distance formula of a point from a plane

(x1-1)2 +(x1+1)2 +(y1-1)2 +(y1+1)2 +(z1-1)2 +(z1+1)2 =10

2x12 + 2y12+ 2z1+ 6 =10

x12 + y12+ z1=2


Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you the answer and detailed
solution very  quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin